OpenStudy (anonymous):

Write the point-slope form of an equation for the line passing through the points (-3, 6) and (2, -9).

OpenStudy (anonymous):

\[m=\frac{ \Delta(y) }{ \Delta(x) }=\frac{ -9-(6) }{ 2-(-3)}=\frac{ -15 }{ 5 }=-3\] \[(y-y1)=m(x-x1)\]\[=y-6=-3(x-(-3))=-3x-9\] y=-3x-9+6 y=-3x-3

OpenStudy (anonymous):

@picklejensen99 get it?

OpenStudy (anonymous):

yes fazeli's right. you need first to get the slope and then you can substitute the points and the slope (M) to the 2nd equation that fazeli gave to you